if a car increases its velocity from 0m/s to 10m/s in 5 seconds,what is the acceleration of the car.
1
Expert's answer
2020-03-25T11:00:39-0400

The acceleration of the car


"a=\frac{v_f-v_i}{t}=\frac{10-0}{5}=2\:\rm m/s^2."